# Understanding Emotion in Discourse: Recognition Insights and Linguistic Patterns for Generation

### v1

@@ Abstract
While Emotion Recognition in Conversation (ERC) has achieved high accuracy, two critical gaps remain: a limited understanding of which architectural choices actually matter, and a lack of linguistic analysis connecting recognition to generation. We address both gaps through a systematic analysis of the IEMOCAP dataset.
For recognition, we conduct a rigorous ablation study with 10-seed evaluation and report three key findings. First, conversational context is paramount, with performance saturating rapidly—90% of the total gain achieved within just the most recent 10–30 preceding turns (depending on the label set). Second, hierarchical sentence representations help at utterance-level, but this benefit disappears once conversational context is provided, suggesting that context subsumes intra-utterance structure. Third, external affective lexicons (SenticNet) provide no gain, indicating that pre-trained encoders already capture necessary emotional semantics. With simple architectures using strictly causal context, we achieve 82.69% (4-way) and 67.07% (6-way) weighted F1, outperforming prior text-only methods including those using bidirectional context.
For linguistic analysis, we analyze 5,286 discourse marker occurrences and find a significant association between emotion and marker positioning ($p<.0001$). Notably, sad utterances exhibit reduced left-periphery marker usage (21.9%) compared to other emotions (28–32%), consistent with theories linking left-periphery markers to active discourse management. This connects to our recognition finding that sadness benefits most from context (+22%p): lacking explicit pragmatic signals, sad utterances require conversational history for disambiguation.
@@ 1 Introduction
Emotion recognition in conversation (ERC) is a central challenge for building socially intelligent dialogue systems, mental health support tools, and empathetic AI agents. Unlike sentence-level emotion detection, ERC requires models to account for conversational context, speaker roles, and subtle pragmatic cues. Recent advances with large pretrained encoders have achieved impressive accuracy—state-of-the-art text-only methods reach 81.4% (4-way) and 64.4% (6-way) weighted F1 on IEMOCAP (dutta2024hcam). Yet these empirical gains come at the cost of transparency: we have little understanding of what architectural choices actually drive performance.
Gap 1: What actually matters?
The field has pursued increasingly complex architectures: hierarchical attention mechanisms (majumder2019dialoguernn), knowledge graph integration (zhong2020ket), external lexicon fusion (tu2022context), and elaborate context modeling (ghosal2019dialoguegcn). Yet we lack systematic understanding of which components genuinely contribute. Most prior studies report single-seed results without statistical tests, making it difficult to distinguish genuine improvements from random variation. Do we really need hierarchical sentence representations? Does external affective knowledge help, or do pretrained encoders already capture this information? How much conversational context is sufficient?
Gap 2: Recognition does not inform generation.
High classification accuracy does not translate to understanding how emotions manifest linguistically. Current ERC models provide no guidance for generating emotionally appropriate text. What linguistic patterns distinguish sad from angry utterances? The role of discourse markers in expressing subjectivity and intersubjectivity is well-established in linguistics (schiffrin1987discourse; beeching2014discourse), yet their positional patterns in ERC remain unexplored. Bridging this gap could benefit both recognition (through linguistically-motivated features) and generation (through actionable production guidelines).
@@ 1.1 Research Questions
We address these gaps through systematic empirical analysis on IEMOCAP, combining rigorous ablation studies with large-scale discourse analysis. Our investigation is guided by four research questions:
@@
    Does conversational context improve emotion recognition, and how much is sufficient?
@@
    Does hierarchical sentence representation help?
@@
    Does external affective lexicon (SenticNet) help?
@@
    Are there emotion-specific discourse marker patterns that could inform generation?
@@ 1.2 Contributions
    We conduct the first ERC study with 10-seed evaluation, paired $t$-tests, and Bonferroni correction, establishing a standard for reliable comparison.
@@
    We isolate the contributions of context, hierarchical structure, and external lexicons, revealing that context is paramount while lexicons provide no benefit.
@@
    We show that different emotions require different amounts of context—Sad benefits most (+22%p) while Angry benefits least (+8%p)—challenging arousal-based explanations.
@@
    We conduct the first large-scale analysis of discourse marker positioning in ERC (5,286 occurrences), revealing that Sad utterances show significantly reduced left-periphery usage.
@@
    We achieve 82.69% (4-way) and 67.07% (6-way) using strictly causal context, outperforming bidirectional methods and enabling real-time deployment.
@@ 2.1 Emotion Recognition in Conversation
Early ERC methods focused on capturing conversational dynamics through recurrent architectures. DialogueRNN (majumder2019dialoguernn) models speaker states across turns, achieving 76.2% on IEMOCAP 4-way classification. COSMIC (ghosal2020cosmic) incorporates commonsense knowledge for context enhancement (77.4%). Recent work has pursued increasingly complex architectures: graph-based models (ghosal2019dialoguegcn), hierarchical attention (ma2022multi), and multimodal transformers (hu2022unimse).
Among text-only methods, HCAM (dutta2024hcam) achieves 81.4% (4-way) and 64.4% (6-way) through hierarchical Bi-GRU with self-attention, while EmoCaps (li2022emocaps) reaches 69.49% on 6-way using emotion capsule networks. Critically, both methods exploit bidirectional context—future utterances unavailable in real-time settings. Our strictly causal approach achieves 82.69% on 4-way, surpassing all text-only methods including bidirectional ones, and 67.07% on 6-way, outperforming six bidirectional methods despite using only past context.
A critical limitation across these methods is their arbitrary context modeling: most fix KK=3 preceding utterances or process entire conversations without considering that different emotions may require different context lengths. While zhang2023variable introduced adaptive instance-level context selection, they still treat all emotions uniformly. We address this through systematic context analysis, exploring whether different emotions exhibit distinct saturation patterns.
@@ 2.2 Discourse Markers and Affective Lexicons
Linguistic theory has long recognized discourse markers as signals of speaker stance and emotion (schiffrin1987discourse). fraser1999discourse classified pragmatic markers by function, while beeching2014discourse demonstrated their tendency to cluster at utterance peripheries. The left periphery typically hosts subjective markers expressing speaker attitude: well signals hesitation, oh marks surprise or realization, and actually introduces contrast or correction (traugott2010intersubjectivity).
Despite extensive theoretical study, computational models have largely ignored these positional patterns. Prior ERC models treat all token positions equally, missing potential emotional cues encoded in discourse structure. We address this gap through systematic analysis of discourse markers in IEMOCAP, examining whether positional patterns vary by emotion. To our knowledge, this is the first large-scale corpus study linking discourse marker positions to emotion categories in conversational speech.
Affective lexicons provide complementary emotional knowledge. While early resources like WordNet-Affect (strapparava2004wordnet) offered categorical labels, SenticNet (cambria2024senticnet) maps concepts onto psychological dimensions. Knowledge-enhanced models (zhong2020ket; tu2022senticgat) show improvements, but whether lexicons add unique information beyond contextual encoders remains unclear—a question we address empirically.
@@ 3.1 Task and Dataset
We conduct experiments on the IEMOCAP dataset (busso2008iemocap), which contains approximately 12 hours of audiovisual data from 10 actors performing scripted and improvised dyadic conversations across 110 dialogue sessions. The dataset exhibits substantial variability in dialogue length (Figure 1a), with sessions ranging from 24 to 267 utterances (mean: 91.7, median: 67.5). Approximately 70% of dialogues contain fewer than 100 utterances, while the 95th percentile reaches 196 utterances, highlighting the long-tail distribution inherent in conversational data.
@@
We define the context length $K$ as the number of preceding turns included as strictly past-only conversational history for a target utterance. Because there is no established stopping rule for how much prior dialogue is sufficient in ERC, we perform an exhaustive bottom-up sweep $K\in\{0,1,\ldots,K_{\max}\}$ to characterize the performance–context trade-off and identify saturation behavior. We set $K_{\max}$ to the maximum dialogue length (in turns) observed in the corpus distribution used to construct dialogue histories. Importantly, this sweep is used strictly for analysis (saturation characterization) and not for test-set model selection; we do not choose $K$ to maximize test performance.
For a fair comparison with state-of-the-art models (dutta2024hcam), we follow the standard speaker-disjoint split strategy: Sessions 2–4 serve as training data (6,072 utterances from 68 dialogues), Session 1 as validation (1,819 utterances from 20 dialogues), and Session 5 as test set (2,196 utterances from 22 dialogues). This split ensures speaker independence between training and evaluation, with each dialogue containing an average of 89–100 utterances across splits. The sentence-level structure within utterances (Figure 1b) shows that most utterances contain 2–5 sentences, motivating our comparison of flat versus hierarchical encoding strategies.
We evaluate on two emotion taxonomies commonly used in the literature. The 4-way classification task considers angry (1,103 utterances, 19.9%), happy (1,636, 29.6%), sad (1,084, 19.6%), and neutral (1,708, 30.9%) emotions, where the excited category is merged into happy (Figure 1c). The 6-way task extends this to include excited (1,041, 14.1%) and frustrated (1,849, 25.1%) as separate categories (Figure 1d). We use weighted F1-score as our primary metric to account for class imbalance inherent in conversational emotion data. We supplement this with class-wise F1 scores and confusion matrices for detailed error analysis.
@@ 3.2 Discourse Marker Analysis
Beyond recognition accuracy, we conduct a linguistic analysis of discourse markers (DMs) to examine emotion-specific pragmatic patterns that may inform future work on emotion-conditioned dialogue generation. DMs are lexical expressions that signal relationships between discourse segments rather than contributing to propositional content (schiffrin1987discourse; fraser1999discourse). Drawing from established taxonomies (schiffrin1987discourse; fraser1999discourse; traugott2010intersubjectivity; beeching2014discourse), we identify 20 markers occurring in IEMOCAP, including turn-management markers (well, oh), connectives (and, but, so), and stance markers (I think, I guess, maybe, you know, I mean). See Appendix A for the full inventory and frequency distribution.
For each marker occurrence, we record its relative position within the utterance (normalized to $[0,1]$), its periphery classification, and the emotion label of the containing utterance. In discourse, the left periphery (LP) and right periphery (RP) serve asymmetric functions (for a comprehensive review, see beeching2014discourse): LP is where speakers claim the floor and manage topic structure, serving textual and subjective functions, while RP is oriented toward the hearer, serving intersubjective and modalising functions. We operationally define LP as position <0.15<0.15, RP as position >0.85>0.85, and medial otherwise. For consistency with our pooling notation, wmean_pos_rev emphasizes utterance-initial (left-peripheral) tokens, whereas wmean_pos emphasizes utterance-final (right-peripheral) tokens; mean pooling treats all positions uniformly.
To test for emotion-specific positional patterns, we employ ANOVA to compare mean positions across emotions, $\chi^{2}$ tests with Cramér’s $V$ to assess association between periphery categories and emotions, mixed-effects models with dialogue as random intercept to control for dialogue-level variation, and post-hoc pairwise comparisons with Bonferroni correction.
@@ 3.3 Model Architecture
To investigate how emotional information is encoded and propagated in dialogue, we develop two encoder variants: a flat encoder and a hierarchical encoder. The flat encoder processes each utterance as a single sequence, while the hierarchical encoder first encodes individual sentences within an utterance, then aggregates them to form the utterance representation. This allows us to examine whether intra-utterance structure provides additional benefit for emotion recognition.
During embedding generation, we include all utterances from each dialogue, even those without target emotion labels (i.e., labels other than happy, sad, angry, neutral for 4-way classification) or unlabeled utterances. This captures long-range contextual dependencies while excluding these utterances from classification training and evaluation. Utterances within each dialogue are indexed to preserve sequential order for turn-level context modeling. Figure 2 illustrates our architecture.
@@
We compare two encoding strategies: (1) flat encoding, which treats each utterance (turn) as a single sequence and extracts the pooled representation, and (2) hierarchical encoding, which first encodes individual sentences within an utterance, then aggregates them into an utterance-level representation.
We evaluate three pre-trained encoders: BERT-base-uncased (devlin2019bert), RoBERTa-base (liu2019roberta), and Sentence-RoBERTa (NLI-RoBERTa-base-v2) (reimers-2019-sentence-bert). For layer selection, we compare two strategies: (1) avg_last4, which averages the last four transformer layers (layers 9–12), capturing high-level semantic features while avoiding overspecialization of the final layer, and (2) last, which uses only the final layer output.
@@
For utterance-level classification (K=0), the utterance representation is passed through a two-layer MLP with ReLU activation and dropout. When incorporating turn-level context (K>0), we process the sequence of utterance representations through a single-layer unidirectional LSTM and use the final hidden state for classification.
@@
To examine whether external affective knowledge benefits emotion recognition, we integrate SenticNet 7 (cambria2024senticnet), which provides four-dimensional affective ratings (pleasantness, attention, sensitivity, aptitude) for words and phrases. For each utterance, we extract SenticNet features by matching constituent words, aggregate them via mean pooling, and concatenate with the encoder representation. This allows us to test whether pre-trained encoders already capture affective semantics or whether explicit lexical knowledge provides additional benefit.

### v2

@@ Abstract
Despite strong recent progress in Emotion Recognition in Conversation (ERC), two gaps remain: we still lack a clear understanding of which modeling choices materially affect performance, and we have limited linguistic analysis that links recognition findings to actionable cues for generation. We address both gaps via a systematic study on IEMOCAP.
For recognition, we conduct controlled ablations with 10 random seeds and paired tests over seeds (with correction for multiple comparisons), yielding three findings. First, conversational context is the dominant factor: performance saturates quickly, with roughly 90% of the gain observed within our context sweep achieved using only the most recent 10–30 preceding turns (depending on the label set). Second, hierarchical sentence representations improve utterance-only recognition (K=0K{=}0), but the benefit vanishes once turn-level context is available, suggesting that conversational history subsumes much of the intra-utterance structure. Third, a simple integration of an external affective lexicon (SenticNet) does not improve results, consistent with pretrained encoders already capturing much of the affective signal needed for ERC. Under a strictly causal (past-only) setting, our simple models attain strong performance (82.69% 4-way; 67.07% 6-way weighted F1), indicating that competitive accuracy is achievable without access to future turns.
For linguistic analysis, we examine 5,286 discourse-marker occurrences and find a reliable association between emotion and marker position within the utterance ($p<0.0001$). In particular, Sad utterances show reduced left-periphery marker usage (21.9%) relative to other emotions (28–32%), aligning with accounts that link left-periphery markers to active discourse management. This pattern is consistent with our recognition results, where Sad benefits most from conversational context (+22%p), suggesting that sadness often relies more on discourse history than on overt pragmatic signaling in the utterance itself.
@@ 1 Introduction
Emotion Recognition in Conversation (ERC) is a key capability for socially intelligent dialogue systems, mental-health support tools, and empathetic conversational agents. Unlike sentence-level emotion classification, ERC requires models to interpret an utterance in relation to conversational history, speaker dynamics, and pragmatic signals that are often subtle or indirect. Recent text-only approaches built on large pretrained encoders have achieved strong performance on benchmarks such as IEMOCAP (Dutta and Ganapathy, 2024), but progress has been accompanied by a growing mismatch between empirical gains and interpretability: it remains unclear which architectural ingredients actually matter and what linguistic evidence these models are exploiting.
Gap 1: Which modeling choices materially matter in modern ERC?
The ERC literature has introduced increasingly elaborate components—hierarchical encoders (Majumder et al., 2019), graph/knowledge-based modules (Zhong et al., 2020; Ghosal et al., 2019), lexicon fusion (Tu et al., 2022), and sophisticated context modeling (Ghosal et al., 2019). However, many reported improvements rely on single-seed runs and heterogeneous experimental setups, making it difficult to separate robust effects from variance or configuration-specific artifacts. As a result, foundational design questions remain unresolved. How much conversational context is truly necessary before gains saturate? Does modeling intra-utterance sentence structure still help once turn-level context is available? Do external affective lexicons add signal beyond what pretrained encoders already capture?
Gap 2: Recognition rarely translates into linguistic guidance for generation.
High recognition accuracy does not automatically reveal how emotions are expressed in discourse, nor does it provide actionable guidance for emotion-conditioned generation. In linguistics, discourse markers are well known to encode stance, subjectivity, and intersubjectivity (Schiffrin, 1987; Beeching and Detges, 2014). Their distribution at the left and right peripheries of utterances is linked to discourse management and listener-oriented effects (Fraser, 1999; Traugott, 2010; Beeching and Detges, 2014). Yet ERC research has rarely examined whether such positional patterns vary systematically by emotion. Bridging recognition and linguistic analysis could benefit both directions: it can provide interpretable hypotheses about what models exploit, and it can yield concrete production cues for generation.
@@ 1.1 Research Questions
We address these gaps through a systematic study on IEMOCAP that combines controlled ablation experiments with large-scale discourse-marker analysis. Our investigation is guided by four research questions:
@@
    Does conversational context improve emotion recognition, and how much context is sufficient before performance saturates?
@@
    Does hierarchical sentence representation help beyond flat utterance encoding, and under what context conditions?
@@
    Does incorporating an external affective lexicon (SenticNet) help under a simple integration scheme?
@@
    Are there emotion-specific discourse-marker positioning patterns that could inform emotion-conditioned generation?
@@ 1.2 Contributions
    We provide a statistically grounded ablation framework for ERC by reporting 10-seed evaluations with paired significance testing and correction for multiple comparisons, enabling reliable component-level conclusions.
@@
    We isolate the relative contributions of context length, hierarchical structure, and lexicon fusion (Tu et al., 2022), showing that conversational context dominates and saturates quickly, while hierarchical encoding primarily benefits utterance-only settings.
@@
    We show that emotions differ substantially in their dependence on conversational history, with sad benefiting more from context than angry, suggesting that context effects are not fully explained by arousal alone.
@@
    We connect recognition to linguistic evidence by analyzing discourse-marker usage and positioning (Schiffrin, 1987; Fraser, 1999; Beeching and Detges, 2014; Traugott, 2010), demonstrating a reliable association between emotion and marker position, including reduced left-periphery usage in sad utterances.
@@
    Under strictly causal (past-only) constraints, our simple models achieve strong performance on both 4-way and 6-way classification, illustrating that competitive accuracy is attainable without access to future turns—a practical consideration for real-time deployment.
@@ 2.1 Emotion Recognition in Conversation
Early ERC methods focused on capturing conversational dynamics through recurrent architectures. DialogueRNN (Majumder et al., 2019) models speaker states across turns, achieving 76.2% on IEMOCAP 4-way classification. COSMIC (Ghosal et al., 2020) incorporates commonsense knowledge for context enhancement (77.4%). Recent work has pursued increasingly complex architectures, including graph-based models (Ghosal et al., 2019), hierarchical attention (Ma et al., 2022), and multimodal transformers (Hu et al., 2022).
A central axis of variation across ERC systems is how conversational context is operationalized. Several approaches adopt heuristic context windows by concatenating neighboring utterances as raw text; for example, UniMSE concatenates the current utterance with a symmetric window of two preceding and two following turns, which implicitly relies on future context (Hu et al., 2022). Other models assume a fixed number of preceding turns as the local interaction neighborhood (e.g., t=3t{=}3) (Mai et al., 2019), treating context length as a default hyperparameter rather than an empirically justified operating point. For transformer baselines, long contexts are often handled by concatenation with truncation: when the input exceeds the model’s length limit, remote utterances are discarded, imposing an implicit cutoff on usable history (e.g., Shen et al. (2021)).
Among high-performing text-only methods, HCAM (Dutta and Ganapathy, 2024) explicitly introduces inter-utterance context via a bidirectional GRU (Bi-GRU) with self-attention, aiming to leverage information from the full conversation. Their formulation uses Bi-GRU to incorporate contextual signals and adds self-attention to better handle long-range dependencies across dialogues with many utterances. While effective, such bidirectional context access is not available in real-time settings where future turns are unknown.
In contrast, we focus on strictly causal (past-only) modeling and explicitly characterize the performance–context trade-off by sweeping the number of preceding turns and quantifying saturation behavior. This complements prior work by replacing heuristic context choices (fixed windows or implicit truncation) with a systematic analysis of how much past text is sufficient, and whether context requirements differ across emotions.
@@ 2.2 Discourse Markers and Affective Lexicons
Linguistic theory has long recognized discourse markers as signals of speaker stance and emotion (Schiffrin, 1987). Fraser (1999) classified pragmatic markers by function, while Beeching and Detges (2014) demonstrated their tendency to cluster at utterance peripheries. The left periphery often hosts markers expressing speaker stance or discourse management (e.g., well signaling hesitation, oh marking surprise or realization, and actually introducing contrast or correction), while right-peripheral markers can be more hearer-oriented (Traugott, 2010; Beeching and Detges, 2014).
Despite extensive theoretical study, computational ERC models have rarely examined whether such positional patterns vary systematically by emotion. We provide a large-scale corpus analysis of discourse-marker positioning in IEMOCAP and evaluate its association with emotion labels, complementing prior linguistic accounts of peripheral distributions.
Affective lexicons provide complementary emotional knowledge. While early resources such as WordNet-Affect (Strapparava, 2004) offered categorical labels, SenticNet (Cambria et al., 2024) maps concepts onto psychological dimensions. Knowledge-enhanced models (Zhong et al., 2020; Tu and others, 2022) report gains, but whether lexicons add unique information beyond contextual encoders remains unclear—a question we examine empirically under a simple integration scheme.
@@ 3.1 Task and Dataset
We conduct experiments on the IEMOCAP dataset (Busso et al., 2008), which contains approximately 12 hours of audiovisual data from 10 actors performing scripted and improvised dyadic conversations across 110 dialogue sessions. The dataset exhibits substantial variability in dialogue length (Figure 1a), with sessions ranging from 24 to 267 utterances (mean: 91.7, median: 67.5). Approximately 70% of dialogues contain fewer than 100 utterances, while the 95th percentile reaches 196 utterances, highlighting the long-tail distribution inherent in conversational data.
@@
We define the context length $K$ as the number of preceding turns included as strictly past-only conversational history for a target utterance. Because there is no established stopping rule for how much prior dialogue is sufficient in ERC, we perform a bottom-up sweep $K\in\{0,1,\ldots,K_{\max}\}$ to characterize the performance–context trade-off and identify saturation behavior. We set $K_{\max}=200$ based on the dialogue-length distribution (covering the long tail while keeping the sweep computationally tractable). [^1] The maximum dialogue length in IEMOCAP is 267 turns; our choice of $K_{\max}=200$ is driven by the empirical distribution (Figure 1a), where 200 exceeds the 95th percentile, and is used for analysis rather than model selection. Crucially, the sweep is used strictly for analysis (saturation characterization) and not for test-set model selection; we do not choose $K$ to maximize test performance.
When a target utterance has fewer than $K$ preceding turns (e.g., early in a dialogue), we use all available history (a variable-length prefix). For turn-level models, sequences are left-padded for batching and padded positions are masked so they do not contribute to contextual aggregation.
For a fair comparison with state-of-the-art models (Dutta and Ganapathy, 2024), we follow the standard speaker-disjoint split strategy: Sessions 2–4 serve as training data (6,072 utterances from 68 dialogues), Session 1 as validation (1,819 utterances from 20 dialogues), and Session 5 as test set (2,196 utterances from 22 dialogues). This split ensures speaker independence between training and evaluation, with each dialogue containing an average of 89–100 utterances across splits. The sentence-level structure within utterances (Figure 1b) shows that most utterances contain 2–5 sentences, motivating our comparison of flat versus hierarchical encoding strategies. Sentence segmentation for hierarchical encoding is performed with spaCy’s sentencizer after lowercasing and standard token normalization. [^2] We follow common ERC preprocessing practice (lowercasing and tokenization) and use a deterministic sentence boundary detector to ensure reproducibility.
We evaluate on two emotion taxonomies commonly used in the literature. The 4-way classification task considers angry (1,103 utterances, 19.9%), happy (1,636, 29.6%), sad (1,084, 19.6%), and neutral (1,708, 30.9%) emotions, where the excited category is merged into happy (Figure 1c). The 6-way task extends this to include excited (1,041, 14.1%) and frustrated (1,849, 25.1%) as separate categories (Figure 1d). We use weighted F1-score as our primary metric to account for class imbalance inherent in conversational emotion data, and we also report class-wise F1 scores and confusion matrices for error analysis.
@@ 3.2 Discourse Marker Analysis
Beyond recognition accuracy, we conduct a linguistic analysis of discourse markers (DMs) to examine emotion-specific pragmatic patterns that may inform future work on emotion-conditioned dialogue generation. DMs are lexical expressions that signal relationships between discourse segments rather than contributing to propositional content (Schiffrin, 1987; Fraser, 1999). Drawing from established taxonomies (Schiffrin, 1987; Fraser, 1999; Traugott, 2010; Beeching and Detges, 2014), we identify 20 markers occurring in IEMOCAP, including turn-management markers (well, oh), connectives (and, but, so), and stance markers (I think, I guess, maybe, you know, I mean). See Appendix A for the full inventory and frequency distribution.
For each marker occurrence, we record its relative position within the utterance (normalized to $[0,1]$), its periphery classification, and the emotion label of the containing utterance. Following standard discourse accounts of utterance peripheries (Beeching and Detges, 2014), we operationally define the left periphery (LP) as position <0.15<0.15, the right periphery (RP) as position >0.85>0.85, and medial otherwise. For consistency with our pooling notation, wmean_pos_rev emphasizes utterance-initial (left-peripheral) tokens, whereas wmean_pos emphasizes utterance-final (right-peripheral) tokens; mean pooling treats all positions uniformly.
To test for emotion-specific positional patterns, we employ (i) ANOVA to compare mean positions across emotions, (ii) $\chi^{2}$ tests with Cramér’s $V$ to assess association between periphery categories and emotions, and (iii) mixed-effects models with dialogue as a random intercept to control for dialogue-level variation (e.g., $\text{pos}\sim\text{emotion}+(1|\text{dialogue})$). We apply post-hoc pairwise comparisons with Bonferroni correction where appropriate.
@@ 3.3 Model Architecture
To investigate how emotional information is encoded and propagated in dialogue, we develop two encoder variants: a flat encoder and a hierarchical encoder. The flat encoder processes each utterance as a single sequence, while the hierarchical encoder first encodes individual sentences within an utterance, then aggregates them to form the utterance representation. Figure 2 illustrates our architecture.
@@
We compare (1) flat encoding, which treats each utterance as a single sequence and extracts a pooled representation, and (2) hierarchical encoding, which encodes each sentence and then aggregates sentence representations into an utterance vector. We evaluate three pre-trained encoders: BERT-base-uncased (Devlin et al., 2019), RoBERTa-base (Liu et al., 2019), and Sentence-RoBERTa (NLI-RoBERTa-base-v2) (Reimers and Gurevych, 2019). Unless stated otherwise, encoders are used as fixed feature extractors and utterance embeddings are precomputed to isolate the effects of contextual modeling and pooling choices. [^3] This design avoids confounding the analysis with end-to-end fine-tuning instability across seeds.
We compare two layer strategies: (i) avg_last4, averaging the last four transformer layers, and (ii) last, using only the final layer output. We use a maximum input length of 128 subword tokens for the utterance encoder; longer utterances are truncated from the left/right as in standard transformer preprocessing.
Context construction and label availability.
During embedding generation, we include all turns in each dialogue in chronological order, including utterances without target emotion labels or outside the target label set. These turns may appear in the contextual history, but only target-labeled utterances contribute to the training loss and evaluation metrics. This allows long-range history to be represented without altering the supervised label space.
@@
For utterance-level classification (K=0K{=}0), the utterance representation is passed through a two-layer MLP with ReLU activation and dropout. When incorporating turn-level context (K>0K{>}0), we process the sequence of utterance representations via a single-layer unidirectional LSTM and use the final hidden state for classification.
@@
To examine whether external affective knowledge benefits emotion recognition, we integrate SenticNet 7 (Cambria et al., 2024), which provides four-dimensional affective ratings (pleasantness, attention, sensitivity, aptitude) for words and phrases. We lowercase and tokenize each utterance, match SenticNet entries at the token level (and, where available, multiword expressions), and aggregate matched vectors via mean pooling; unmatched tokens contribute zeros. The resulting 4D vector is concatenated with the encoder representation.